OpenStudy (anonymous):

One number is five more than another, and their sum is three less than three times the smaller. Find the numbers. If x represents the smaller number, which equation could be used to solve for x? A) x 2 + 5 = 3x - 3 B) 2x + 5 = 3x - 3 C) 2x + 5 = 3(x - 3)

OpenStudy (anonymous):

@gorv

OpenStudy (gorv):

let a number x another=x+5 sum=x+x+5<3x-3 2x+5<3x-3 5+3<3x-2x x>8

OpenStudy (gorv):

B